OpenStudy (anonymous):

If the half-life of strontium-90 is 29 years, what fraction of strontium-90 absorbed in 1960 remained in people's bones in 1990?

OpenStudy (alekos):

that would be very close to one half

OpenStudy (anonymous):

Right. I need to use\[Q=Q _{0}a ^{t}\]I got the decay rate at 2%, but can't figure out how to get an exact answer.

OpenStudy (alekos):

I get 48%

OpenStudy (alekos):

Formula is Q=Qo*e^-rt where r=-ln2/29

OpenStudy (alekos):

I worked out r using Q/Qo = 1/2 for half life

OpenStudy (alekos):

from here you just substitute r and t=30

OpenStudy (alekos):

and you get 48.8%

OpenStudy (anonymous):

correct answer is 46.85% ? @wrstlr3232 ..I'm not sure

OpenStudy (dbzfan836):

It must be just above 50%

OpenStudy (anonymous):

I got 48.8%. That makes sense. Thanks everyone!